# Benchmarking General-Purpose In-Context Learning

### v5

@@ Abstract
In-context learning (ICL) empowers generative models to address new tasks effectively and efficiently on the fly, without relying on any artificially crafted optimization techniques. In this paper, we study extending ICL to address a broader range of tasks with an extended learning horizon and higher improvement potential, namely General-Purpose In-Context Learning (GPICL). To this end, we introduce two lightweight benchmarks specifically crafted to train and evaluate GPICL functionalities. Each benchmark encompasses a vast number of tasks characterized by significant task variance, facilitating meta-training that minimizes inductive bias. These tasks are also crafted to promote long-horizon in-context learning through continuous generation and interaction. These characteristics necessitate the models to leverage contexts and history interactions to enhance their capabilities, across domains such as language modeling, decision-making, and world modeling. Our experiments on the baseline models demonstrate that meta-training with minimal inductive bias and ICL from the ground up is feasible across all the domains we’ve discussed. Additionally, our findings indicate that the scale of parameters alone may not be crucial for ICL or GPICL, suggesting alternative approaches such as increasing the scale of contexts and memory states.
00footnotetext: Works done at Baidu Inc. and University of Science and Technology of China.00footnotetext: Corresponds with: Fan Wang (fanwang.px@gmail.com) and Yang Cao (forrest@ustc.edu.cn)00footnotetext: Source code available at https://github.com/FutureAGI/L3C
@@ 1 Introduction
The success of large-scale generative language models can be primarily attributed to two key factors: 1. Zero-shot generalization derived from the extensive accumulation and storage of knowledge within the model parameters during the pre-training and fine-tuning stages (Kojima et al., 2022); 2. In-context learning(ICL) to distill knowledge from the context during the inference stage (Brown et al., 2020). The ability of zero-shot generalization remains unchanged post-training, which parallel the concept of “innate abilities” (Koulakov et al., 2021) of biology. The ICL capability allows the agent to handle various unseen tasks without modifying the model parameters, only resulting in changes to the hidden states and memories. Currently, the ICL capabilities of large language models are confined to tasks that solely involve interpreting natural language tokens and following instructions (i.e., mimicking demonstrations). However, they generally lack the ability to continually interact with real-world environments and engage in reinforcement learning from contexts. Additionally, extremely complex tasks that require substantial amounts of context still pose significant challenges.
@@
Figure 1: Conceptual diagram of in-context learning (ICL) and general-purpose in-context learning (GPICL)
There are significant gap between artificial intelligence and biological intelligence, particularly regarding the functionality to continuously learn throughout their lifespans (Parisi et al., 2019). The enormous lifelong learning and adaptation potential of biological neural networks, forming abrupt contrast with their limited innate abilities at the initial stages (e.g., the stage of mammalian infancy), are also used to guide the design of machine intelligence systems (Zador, 2019; Wang et al., 2022; Schmidgall et al., 2024). A promising area in this line of research is general-purpose in-context learning (GPICL) (Kirsch et al., 2022; 2023). GPICL involves meta-training across a wide range of task classes and learning to interpret context within a variety of highly diverse tasks. However, the characteristic of diversity in this context can be challenging to quantify. We aim to enrich the concept of GPICL by incorporating additional features, as shown in Figure 1. Consider an agent that ideally utilizes all available data, including meta-training data and contextual information; The GPICL system should corresponds to low zero-shot capability, high ICL potential, and an extended ICL horizon.
@@
Importance of GPICL. The objective of extending the in-context learning (ICL) horizon and potential is widely acknowledged. However, suppressing zero-shot capability seem to contradict the prevailing trend in large language models. To elucidate its importance further, we emphasize three key points:
First, theoretically, the lower zero-shot capability in meta-training indicates a reduced reliance on inductive biases, thereby demonstrating sufficient diversity of meta-training tasks. As a result, the meta-training process focus primarily on the functionality of ICL instead of task-specific knowledge. Conversely, the attainment of zero-shot capability might suggest that the tasks possess a common knowledge base. While it could lead to faster adaption within the scope of these tasks, it might potentially restrict the model’s applicability beyond the scope of these tasks.
Second, practically, meta-training the learning functionality requires an enormous number of tasks, which is challenging to collect in the real world. The use of abundant synthetic data in pre-training and meta-training can be essential, but tend to be non-factual and biased, presenting challenges such as the sim-to-real gap. Ideally, the meta-training stage of GPICL results in the loss of much task-specific knowledge while preserving the ICL functionality, thus helps prevent the model from the biases in the synthetic datasets. Building on this point, an alternative approach could be developed to replace the current mainstream learning pipeline of Pre-training & Fine-tuning & ICL with Meta-training & GPICL (Figure 2). In the new pipeline, initially, massive synthesized data, potentially containing a wealth of nonfactual and biased information, is used exclusively to cultivate the GPICL functionality. Subsequently, a considerable amount of high-quality, unbiased, well-aligned, real-world data is employed to enrich the contexts and memory, enabling the agent to achieve capabilities approaching those of humans.
Third, since a model’s performance typically depends on both in-context learning (ICL) and zero-shot generalization capabilities, it is challenging to investigate and improve ICL independently. Isolating the ICL capability alone would be beneficial for analysis on model architectures and learning algorithms.
Guidelines for designing GPICL benchmarks. We note that a major impediment to research in this area is the lack of benchmarks that adhere to the GPICL standard. We suggest that those benchmarks should satisfy the following standard:
  • Vast number of variant tasks with minimal inductive bias: It is desirable to ensure that the meta-training and meta-testing tasks are sufficiently diverse, preventing the transfer of task-specific knowledge among different tasks. This guarantees that the model’s performance is primarily driven by its in-context learning capabilities rather than inductive biases. It is also essential to require numerous tasks to prevent the model from hacking in-context learning (ICL) by merely memorizing all possible tasks and performing task identification, a limitation that many meta-learning benchmarks fall into when given limited tasks, as is discussed in Kirsch et al. (2022).
  • Interactive and generative tasks: GPICL should include task exploration and completion processes that are both interactive and generative, to ensure obtaining complex functionalities such as in-context reinforcement learning. Consequently, benchmarks that solely focus on encoding a single sample characterized by extended contexts and features and conducting single-step predictions or classifications do not meet the criteria.
  • Lifelong in-context learning: To guarantee the capability of adapting complex tasks, GPICL should handle very long-term dependencies that can scale up to millions or even billions of steps, effectively supporting the learning process over extensive periods (ICL horizon $T$ should be “from cradle to grave”). It is important to clarify that $T$ should denote the “minimum steps” possible given “ideal ICL functionality”, excluding scenarios requiring extensive contexts that are primarily due to inefficient learning functionalities.
Contribution of this paper: We propose two critical benchmarks that adhere to the GPICL standard and are capable of generating unlimited synthetic data: 1. Meta-Language: We create a randomized yet consistent language pattern to meta-train models, demonstrating that a model, without prior exposure to natural language corpora, can acquire the ability to learn a new language from scratch through ICL; 2. Maze World: We generate an unlimited number of tasks within randomized mazes, where the model is meta-trained solely through context to learn exploration and navigation. For both benchmarks, we provide limited experiment results of baseline models based on Transformer (Vaswani et al., 2017) backbone and its variants. The primary goal of this paper is neither to introduce a new model or learning methodology, nor to provide an exhaustive survey of the existing ICL systems; rather, we aim to demonstrate through these baselines the potential for these benchmarks to benefit the research of ICL and GPICL. One crucial conclusion from these limited experiments may be that the scale of parameters alone is not crucial for ICL or GPICL. This finding aligns with previous research indicating that the size of contexts, and the scale of the memory states can play a more significant role (Kirsch et al., 2022; Wang et al., 2022) in learning from contexts.
@@ 2.3.1 Towards Meta-Language Model
Figure 3: Perplexity per position on evaluation data sampled by different language complexity.
@@ 2.3.2 Scaling Laws for GPICL
Figure 4: Evaluating language models of various sizes on Meta-Language benchmark with differing levels of complexity ($n$).
@@ 3.3 Baseline Performances
(a) Maze $15\times 15$
(b) Maze $25\times 25$
(c) Maze $35\times 35$
@@
(a) Maze $15\times 15$
(b) Maze $25\times 25$
(c) Maze $35\times 35$
Figure 9: Results regarding the in-context improvement of world modeling in Maze World. We examine the compound error associated with a 9-step auto-regressive future predicting process, using the learned world models that differ in the context length ($t\in\{1,100,1000,2000\}$). The figures presented compare different models to the context-free world model as the baseline. The horizontal axis measures the number of imagined steps, with each curve illustrating a distinct context length. The mean square error is averaged across 64 tasks, and the shaded area provides a 95% confidence interval for these measurements.
@@
In-Context Improvement of World Modeling. Figure 9 explores the performance of world models in interactive evaluations. At steps $1$, $100$, $1000$, and $2000$ of each task, the world model is tasked with “imagining” $9$ steps into the future. For the smaller $15\times 15$ mazes, we observe a significant improvement in world modeling as context length increases from $1$ to $1000$, evidenced by a continuous decrease in prediction error. However, this improvement plateaus between $1000$ and $2000$ steps. With sufficient context, the cumulative error over 9 steps remains lower than that of an average guess, which is the performance of the context-free model. For larger mazes of $25\times 25$ and beyond, and when predicting more than 5 steps into the future, none of the trained models outperform an average guess. Additionally, the partial-context model does not enhance its predictive capabilities as context length increases, which aligns with expectations. The results indicate that the proposed model demonstrates a reasonable capacity to utilize contextual information for future predictions, yet there is considerable room for improvement, especially in solving larger-scale mazes and the efficient utilization of long-term dependencies. For further insights and cases, refer to Appendix A.4.
@@ 5 Conclusions
It is important to note that although we are seeing the potential of application of those datasets in areas including language modeling and indoor navigation, the synthetic data generated by those benchmarks are not yet validated to be ready for a direct help to application. This work can be expanded along two directions: 1. The development of more realistic GPICL benchmarks, for instance, by adding greater complexity to the meta-language; 2. Utilizing the benchmarks to uncover more sophisticated models (including memory-augmented models) and optimization techniques (such as reinforcement learning) to further extend the GPICL horizon and enhance performance gains.
@@ A.1 Model Description
The comprehensive configurations of the baseline models are outlined in Table 1. Notably, within the Maze World framework, each temporal increment is encoded by a pair of tokens: an observation and its corresponding action. For the optimization of the Meta-Language and Maze World baseline models, we employed the Noam decay scheduler [Vaswani et al., 2017]. This scheduler peaks at a learning rate of $10^{-3}$ following an initial warm-up phase spanning $1,000$ steps.
For the image encoder (11M parameters) and image decoder (14M parameters), we employ convolutional and deconvolutional layers augmented with ResNet blocks [He et al., 2016]. The detailed structure of the VAE is depicted in Figure 10. Initially, we train the VAE in isolation by setting $\alpha_{wm}$ and $\alpha_{pm}$ in Equation 8 to 0. Subsequently, we freeze the parameters of the encoder and decoder and train the causal transformer only.
@@
(a) Meta-Language Model
(b) World Modeling
(c) Policy Modeling
(d) World Modeling
(e) Policy Modeling
Figure 12: The training loss with respect to the steps of optimization. Notice that the partial-context and context-free causal transformers are initialized from the trained full-context transformers.
Figure 12 presents the training loss versus the number of iterations for the Meta-Language and Maze World baseline models. During training baselines for Meta-Language tasks, we observed that standard-sized models converge relatively slowly on the training set. To address this, we conducted additional warm-up epochs on a simpler pre-training dataset with $n=2,3$ before transitioning to the main pre-training dataset. During training the baselines for Maze World tasks, we found there was mutual interference between the world modeling loss ($\mathcal{L}_{wm}$) and policy modeling loss ($\mathcal{L}_{pm}$), especially in the initial stages. Our meta-training underwent several phases where we continuously adjusted the loss weights $\alpha_{wm}$ and $\alpha_{pm}$ in Equation 8. To further accelerate and stabilize the convergence of the world model, in addition to these losses, we introduced an auxiliary loss $\mathcal{L}_{z}=\sum_{t}||\hat{z}_{t}-z_{t}||^{2}$ with a weight of $\alpha_{z}=0.05$. During the final stage, the hyperparameters are adjusted to $\alpha_{pm}=0.25$ and $\alpha_{wm}=0.70$. However, we recommend starting with $\alpha_{wm}>0.90$ and $\alpha_{pm}<0.10$ as the policy model training loss is more volatile initially. The context-free and partial-context causal transformers are initialized from the full-context causal transformer, and the hyperparameters remain unchanged.

### v6

@@ Abstract
In-context learning (ICL) empowers generative models to address new tasks effectively and efficiently on the fly, without relying on any artificially crafted optimization techniques. In this paper, we study extending ICL to address a broader range of tasks with an extended learning horizon and higher improvement potential, namely General Purpose In-Context Learning (GPICL). To this end, we introduce two lightweight benchmarks specifically crafted to train and evaluate GPICL functionalities. Each benchmark encompasses a vast number of tasks characterized by significant task variance. These tasks are also crafted to promote long-horizon in-context learning through continuous generation and interaction, covering domains such as language modeling, decision-making, and world modeling. The benchmarks necessitate the models to leverage contexts and history interactions to enhance their capabilities, which we believe to be the key characteristics of GPICL. Our experiments indicate that the diversity of training tasks is positively correlated with the ability to generalize with ICL, but inversely correlated with zero-shot capabilities. Additionally, our findings indicate that the scale of parameters alone may not be crucial for ICL or GPICL, suggesting alternative approaches such as increasing the scale of contexts and memory states.
00footnotetext: source code will be available in the final version
@@ 1 Introduction
The success of large-scale generative language models can be primarily attributed to two key factors: 1. Zero-shot generalization derived from the extensive accumulation and storage of knowledge within the model parameters during the pre-training and fine-tuning stages (Kojima et al., 2022); 2. In-context learning(ICL) to distill knowledge from the context during the inference stage (Brown et al., 2020). The ability of zero-shot generalization remains unchanged post-training, which parallel the concept of “innate abilities” (Koulakov et al., 2021) of biology. The ICL capability allows the agent to handle various unseen tasks without modifying the model parameters, only resulting in changes to the hidden states and memories. Currently, the ICL capabilities of large language models are confined to natural language-based tasks and relatively naive tasks such as following instructions and mimicking demonstrations. However, they generally lack the ability to continually interact with real-world environments and engage in reinforcement learning from contexts. Additionally, extremely complex tasks that require substantial amounts of context still pose significant challenges.
@@
Figure 1: Conceptual diagram of in-context learning (ICL) and general purpose in-context learning (GPICL)
On the other hand, there are significant gap between artificial intelligence and biological intelligence, particularly regarding the functionality to continuously learn throughout their lifespans (Parisi et al., 2019). The enormous lifelong learning and adaptation potential of biological neural networks that forms abrupt contrast with their limited innate abilities at the initial stages (e.g., the stage of mammalian infancy) are also used to inspire the design of machine intelligence systems (Zador, 2019; Wang et al., 2022; Schmidgall et al., 2024). Among those a promising topic is general purpose in-context learning (GPICL) (Kirsch et al., 2022; 2023). GPICL involves meta-training across a wide range of task classes and learning to interpret in context to generalize to new tasks. However, the characteristic of diversity in this context can be challenging to quantify. We aim to enrich the concept of GPICL by incorporating additional features, as shown in Figure 1. We consider the meta-learning framework in which a model is meta-trained in the outer loop to utilize contextual information for task adaptation in the inner loop. We identify a model with GPICL capability by three main characteristics: 1. Low zero-shot performance; 2. High ICL potential; 3. A long ICL horizon.
@@
We propose two lightweight benchmarks that adhere to the GPICL standard and are capable of generating unlimited synthetic data: 1. Meta-Language: This benchmark generates random yet consistent language patterns for meta-training the ICL functionality, enabling it to master a novel language from scratch. 2. Maze World: This benchmark consists of a variety of mazes with different targets that are randomly generated for navigation purposes. It is designed for training and evaluating models capable of navigating in various scenarios through observation, mapping, localization, and planning. We provide experiments of baseline models composed of language modeling, world modeling, and decision modeling based on the Transformer backbone and its variants, as described in Vaswani et al. (2017). It is shown that: 1. With auto-regressive models, both benchmarks exhibit low zero-shot performance and high potential for generalization with ICL; 2. Increasing the diversity of meta-training tasks can naturally lead to decreased zero-shot performance, improved generalization to unseen tasks, and an extended ICL horizon, implying that sufficiently diverse tasks can lead to effective GPICL; 3. The performance of GPICL is not strongly correlated with the scale of parameters once a certain threshold is exceeded. On the other hand, the size of the context significantly influences the model’s performance, which aligns with the findings in Kirsch et al. (2022); Wang et al. (2022). The results strongly suggest that, in evaluations, the position-wise performance of the auto-regressive model should be investigated in detail, rather than averaging the performance across different context lengths.
Our findings reveal a potential alternative approach to replace the current mainstream learning pipeline of large-scale language models, as shown in Figure 2. Typical Large Language Models (LLMs) go through three stages: Pre-training, Fine-tuning, and ICL to train and customize. With GPICL, it is possible to simplify this pipeline to just two stages: Meta-training and GPICL. Moreover, given the data efficiency of GPICL compared with gradient-based methods, it is possible to base meta-training on low-quality synthetic data and use high-quality data more efficiently to master novel skills.
@@ 2.3.1 Towards Meta-Language Model
Figure 3: Comparing evaluations on meta-language sampled by different language complexity. Horizontal axis represent the position or context length, and vertical axis represent the perplexity (the lower the better).
@@ 2.3.2 Scaling Laws for GPICL
Figure 4: Comparing performances of language models of different sizes on the the same test set. Horizontal axis represent the position or context length, and vertical axis represent the perplexity (the lower the better).
@@ 3.3 Baseline Performances
Figure 9: Results regarding the in-context improvement of world modeling in Maze World. We assess the mean square error (MSE) for forecasting subsequent frames. When $k=1$, it indicates the error associated with predicting the immediately following frame. When $k=4$, it denotes the cumulative error for predicting 4 frames ahead in an autoregressive manner. The predictions are initiated from various context lengths (or time step), with $t$ taking values from the set {1, 100, 1000, 2000}. The mean square error is averaged across 64 procedural generated tasks, and the shaded area provides a 95% confidence interval for these measurements.
@@
In-Context Improvement of World Modeling. In Figure 9, the world model is evaluated by comparing the ground truth future frames and predicted future frames at steps $1$, $100$, $1000$, and $2000$ of each interactive run. We demonstrate the performance of world models in predicting both the immediate future ($k=1$) and a slightly more distant future ($k=4$). The latter is more reliant on long-term memory, making it inherently more difficult. In the case of smaller $15\times 15$ mazes, we notice a substantial enhancement in world modeling as the context length expands from $1$ to $1000$, which is validated by a consistent reduction of prediction errors for both $k=1$ and $k=4$. For larger-scale mazes, the challenge of predicting $k=4$ is amplified, as retaining the entire map in memory becomes increasingly challenging. The partial-context model does not enhance its predictive capabilities as context length increases over $100$, which aligns with expectations that it can not keep a long-term memory. The results indicate that the full-context causal transformer demonstrates a reasonable capacity to utilize contextual information for world modeling, yet there is considerable room for improvement, especially in solving larger-scale mazes and the efficient utilization of long-term dependencies. For further insights and cases, refer to Appendix A.4.
4 Emergence of Generalizability
Figure 10: Performance of the meta-language model trained with 500,000 sequences generated from varying numbers of pre-selected tasks. The vertical axis represents the average perplexity in evaluation (lower values indicate better performance), and the horizontal axis represents the context length. Each point in the seen tasks evaluation is the average of 4,000 sequences generated from the pre-selected tasks. Each point in the unseen tasks evaluation is based on the PG-19 test set.
Figure 11: Performance of the causal transformer model trained with 64,000 Maze World sequences generated from varying numbers of pre-defined tasks. The left vertical axis represents the accumulated reward (higher values indicate better performance), while the right vertical axis represents the prediction error of world modeling in the interactive evaluation (lower values indicate better performance). The horizontal axis represents the context length or time steps. For the prediction error of world modeling, we evaluated the mean squared error (MSE) for predicting 1 and 4 steps into the future at context lengths of 1, 100, 1000, and 2000. Each evaluation point is the average of 64 runs, with the shaded area representing the 95% confidence interval.
To illustrate the relationship between the two benchmarks and GPICL, we delve into how the ICL capability is linked to the diversity of meta-training data in these benchmarks. In this section, we pre-select a set of tasks to generate the sequences (dataset) for meta-training. Unlike the procedural generation of tasks for each sequence, this approach reduces the diversity of both the task-set and the dataset, by allowing multiple sequences to be derived from a single task. For example, in Meta-Language, each task is defined by the parameters of the n-gram generator ($\theta$), allowing for the generation of different sequences from a single task through softmax sampling. Similarly, in Maze World, each task is characterized by the layout, textures, and the position of the PNTs (potential navigation targets), with different trajectories being generated by sampling behavior policies. To examine the effect of task diversity on ICL capability, we create datasets of the same size but with varying numbers of pre-defined tasks. In Meta-Language, we generate 500,000 sequences (2 billion tokens) using 10, 1,000, and 100,000 pre-sampled tasks, respectively. In Maze-World, we synthesize 64,000 sequences (126 million frames) with 1, 10, 100, and 10,000 pre-sampled tasks for the $15\times 15$ mazes. Given that the performance of standard-sized models and small-sized models was nearly indistinguishable in our previous experiments, we use only small-sized models for both benchmarks. Each meta-trained model is assessed on both ’seen’ tasks (those used to generate the meta-training data) and ’unseen’ tasks. The outcomes are displayed in Figures 10 and 11.
Several conclusions have been validated across both benchmarks: First, we observe a continuously increasing asymptotic performance (the performance with sufficiently long context) in unseen tasks as the number of meta-training tasks increases, a phenomenon also noted by Kirsch and Schmidhuber (2021). This finding confirms that the capability for task generalization is enhanced. Second, which is less discussed in the previous work, for seen tasks, performance does not consistently improve with the increase in the number of tasks; in contrast, it declines in most of the groups. Additionally, we notice an expansion of the ICL horizon, identified by the length of context during the performance improvement phase, and a reduction in zero-shot performance, identified by the level of performance at the first time step (where context length is 0), in both seen and unseen tasks. We hypothesize that this may be a consequence of the model transitioning from task-identification to learning-to-learn. In this process, the model retains less specific information about each task but gains improved generalizability and the ability to learn through ICL. The decline in performance for seen tasks could also be attributed to the insufficient context length we employ (up to 2,048 and 4,096) and the size of the dataset. Last, we observe a slower convergence in meta-training as the number of tasks increases, despite the dataset size remaining constant (Figure 15). This suggests that the process of learning to learn is more challenging than mastering specific tasks. It is important to note that the observed increase in generalizability, the expansion of the ICL horizon, and the reduction in zero-shot capability are all in line with the definition of GPICL that we previously mentioned.
@@ 6 Conclusions
It is important to note that although we are seeing the potential of application of those datasets in areas including language modeling and indoor navigation, the synthetic data generated by those benchmarks are not yet validated to be ready for a direct help to application. This work can be expanded along two directions: 1. The development of more realistic GPICL benchmarks, for instance, by including additional complexity to the meta-language; 2. Utilizing the benchmarks to uncover more sophisticated models (including memory-augmented models) and optimization techniques (such as reinforcement learning) to further extend the GPICL horizon and enhance performance gains.
@@ A.1 Model Description
The comprehensive configurations of the baseline models are outlined in Table 1. Notably, within the Maze World framework, each temporal increment is encoded by a pair of tokens: an observation and its corresponding action. For the optimization of the Meta-Language and Maze World baseline models, we employed the Noam decay scheduler [49]. This scheduler peaks at a learning rate of $10^{-3}$ following an initial warm-up phase spanning $1,000$ steps.
For the image encoder (11M parameters) and image decoder (14M parameters), we employ convolutional and deconvolutional layers augmented with ResNet blocks [14]. The detailed structure of the VAE is depicted in Figure 12. Initially, we train the VAE in isolation by setting $\alpha_{wm}$ and $\alpha_{pm}$ in Equation 8 to 0. Subsequently, we freeze the parameters of the encoder and decoder and train the causal transformer only.
@@
(a) $-logP$ (Meta-Language-Model)
(b) $\mathcal{L}_{wm}$ (full-context)
(c) $\mathcal{L}_{pm}$ (full-context)
(d) $\mathcal{L}_{wm}$ (partial-context))
(e) $\mathcal{L}_{pm}$ (partial-context)
Figure 14: The training loss with respect to the steps of optimization in training with procedural generated tasks and sequences. Notice that the partial-context and context-free causal transformers are initialized from the trained full-context transformers.
(a) $-logP$ (Small-sized)
(b) $\mathcal{L}_{wm}$ (Small-sized)
(c) $\mathcal{L}_{pm}$ (Small-sized)
Figure 15: The training loss with respect to the steps of optimization in training with sequences generated from pre-selected tasks.
Figure 14 presents the training loss versus the number of iterations for training the baseline in section 2 and A.5. During training baselines for Meta-Language tasks, we observed that standard-sized models converge relatively slowly on the training set. To address this, we conducted additional warm-up epochs on a simpler pre-training dataset with $n=2,3$ before transitioning to the main pre-training dataset. During training the baselines for Maze World tasks, we found there was mutual interference between the world modeling loss ($\mathcal{L}_{wm}$) and policy modeling loss ($\mathcal{L}_{pm}$), especially in the initial stages. Our meta-training underwent several phases where we continuously adjusted the loss weights $\alpha_{wm}$ and $\alpha_{pm}$ in Equation 8. To further accelerate and stabilize the convergence of the world model, in addition to these losses, we introduced an auxiliary loss $\mathcal{L}_{z}=\sum_{t}||\hat{z}_{t}-z_{t}||^{2}$ with a weight of $\alpha_{z}=0.05$. During the final stage, the hyperparameters are adjusted to $\alpha_{pm}=0.25$ and $\alpha_{wm}=0.70$. However, we recommend starting with $\alpha_{wm}>0.90$ and $\alpha_{pm}<0.10$ as the policy model training loss is more volatile initially. The context-free and partial-context causal transformers are initialized from the full-context causal transformer, and the hyperparameters remain unchanged.
Figure 15 presents the training loss versus the number of iterations for the experiments in section 4. Both group are experimented on small-sized models.